A constant net force of 350 N is applied to accelerate a cart in a frictionless surface from rest to a velocity of 65 m/s in 20 s. What is the mass of the cart?


1
Expert's answer
2022-04-11T12:44:32-0400

The acceleration of the cart

"a=\frac{\Delta v}{\Delta t}=\frac{65}{20}=3.25\:\rm m/s^2"

The mass of cart

"m=\frac{F}{a}=\frac{350}{3.25}=108\:\rm kg"